OpenStudy (anonymous):

Number 16 Please! @iheartfood

OpenStudy (anonymous):

Well each triangle has a total of 180 degrees right?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

So Each triangle and numbers can be looked at that way. Look at angle 2 first A line is 180 degress We know that one part of the angle is 125 180 - 125 = 55 Angle 2 is 55

OpenStudy (anonymous):

To find angle 1, we only have to subtract from 180 180 - 55 - 50 = 75

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Is That How You Set It Up?

OpenStudy (anonymous):

Yes You could do it that way. Angle 3 is the exact same as angle 2, being 55 degrees because there is another straight line with an attached 125 degrees

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

To find angle 4: 180 - 55 -85 = 40

OpenStudy (anonymous):

Angle 5 is attached to angle 4 180 - 40 = 140 degrees

OpenStudy (anonymous):

Angle 6 is on the same 180 degree line as 5 180 - 140 = 40

OpenStudy (anonymous):

So It Would Be- 75, 55, 55, 40, 140, 40, 75, 65, 115??????

OpenStudy (anonymous):

Whatever I've been saying yeah. XD That looks right to me

OpenStudy (anonymous):

Alright Thanks.